Given P(A) = 0.10, P(B) = 0.12, P(C) = 0.21, P(A ∩ C) = 0.05, and P(B ∩ C) = 0.03,

solve the following:

a. P(A ∪ C) =

b. P(B ∪ C) =

c. P(A ∪ B) =
Answer #1

a. P(A ∪ C) =P(A)+P(C)=P(A and C)

=0.10+0.21-0.05

=0.26

b. P(B ∪ C) =P(B)+P(C)-P( B and C)

=0.12+0.21-0.03

=0.30